# Rectified Iterative Disparity for Stereo Matching

### v2

@@ Abstract
Frankly speaking, this work serves as a preliminary exploration for SR-Stereo, and the paper writing requires some refinement. However, the incorporation of uncertainty in our approach may provide valuable insights for your research. I highly recommend reading the SR-Stereo paper [27] directly to gain a deeper understanding of the topic.
@@ V-C Estimation of Disparity Uncertainty
TABLE III: Ablation study of disparity rectification. The baseline is IGEV-Stereo. All models are trained on Scene Flow. The final configuration of the UDC and UDR is underlined. Bold: Best.
Methods Variations Scene Flow Middlebury-H ETH3D
EPE(px) $>3px(\%)$ EPE(px) $>2px(\%)$ EPE(px) $>1px(\%)$
Baseline - 0.72 3.65 1.28 8.44 0.492 4.49
$m=2$ 0.69 3.54 1.14 8.05 0.421 4.15
+ UDC $m=3$ 0.70 3.58 1.12 7.55 0.467 4.59
$m=4$ 0.69 3.52 1.32 7.98 0.474 4.54
+ UDR $s=2.0$ 0.71 3.59 1.08 7.69 0.486 4.44
$s=3.0$ 0.71 3.57 1.08 7.37 0.479 4.40
@@ V-D Ablation Study
Configuration Exploration: We explore the optimal configuration of the proposed method. Table III shows the results of UDR and UDC with different configurations. In the vast majority of cases, UDR and UDC outperform the baseline on several datasets. We argue that the disparity distributions on different datasets are very variable, which makes it difficult to generalise the method trained under a fixed range of disparities to another dataset. However, the range of uncertainty is stable. Therefore, UDR and UDC are two beneficial complements, especially for the exploitation of uncertainty information, which significantly improves the generalisation performance of the model. In addition, to demonstrate the generality of the proposed DR loss over different iteration-based methods, we conduct experiments in both RAFT-Stereo and IGEV-Stereo. As shown in Table IX, both methods achieve improvements on $D1_{error}$ for all configurations of DR loss. Notice that when $\beta$ is changed from 0.5 to 0.1, the rectification weights are decreased as a whole, but $D1_{error}$ is optimally improved. This phenomenon demonstrates the focus of DR loss on small error pixels, especially when the weights are more skewed.
TABLE IV: Ablation study of DR loss. All models are trained on Scene Flow. The final configuration of the DR loss is underlined. Bold: Best.
Methods $\alpha$ $\beta$ Scene Flow Middlebury-H ETH3D
EPE(px) $>3px(\%)$ EPE(px) $>2px(\%)$ EPE(px) $>1px(\%)$
RAFT-Stereo (R.) - - 0.79 4.25 1.98 14.67 0.40 5.24
R. + DR loss 0.125 0.1 0.82 3.61 2.21 11.30 0.34 3.61
0.125 0.5 0.77 3.91 1.88 12.51 0.36 4.13
0.250 0.5 0.77 4.04 1.96 13.00 0.32 4.27
IGEV-Stereo (I.) - - 0.72 3.65 1.28 8.44 0.50 4.49
I. + DR loss 0.125 0.1 0.80 3.26 1.20 7.47 0.50 3.99
0.125 0.5 0.71 3.37 1.27 7.79 0.50 4.11
0.250 0.5 0.72 3.56 1.21 7.92 0.48 3.97
TABLE V: Ablation study and complexity of DR-stereo. The baseline is IGEV-Stereo. The last two columns are the results when the size of input image is $1248\times 384$. Bold: Best.
@@
Combinations of proposed methods: As shown in Table V, we experiment with multiple combinations of the proposed methods. Comparing the results with those in Tables III and IX, our methods achieve further improvements after combination. In addition, we record the increase in the number of parameters, time consumption after inserting the design modules. It can be seen that inserting the UDR and UDC puts little burden for the model. As we present in Section IV-B, the combination of UDR and UDC does not increase the number of parameters compared to a single method, since they share parameters from the UEC module.
@@ V-F Benchmarks
In this section, we compare DR-Stereo with the state-of-the-art methods published on Scene Flow and KITTI. Tables VI shows the quantitative results. With similar training strategies, DR-Stereo achieves a new SOTA EPE on the Scene Flow test set. Evaluation results on the KITTI benchmark show that DR-Stereo achieves the best performance on the vast majority of metrics. At the time of writing, our method outperforms all published methods on the online KITTI 2015 leaderboard.
TABLE VI: Quantitative evaluation on Scene Flow and KITTI 2015. Bold: Best.
@@ V-G Zero-shot Generalization
We pre-train DR-Stereo on Scene Flow and test it directly on Middlebury 2014 and ETH3D. As shown in Table VII, Our DR-Stereo achieves very competitive generalisation performance. Compared with the original optimal method IGEV-Stereo, our method achieves an overall improvement.
TABLE VII: Synthetic data generalization experiments. We pre-train our model on Scene Flow and test it directly on Middlebury 2014 and ETH3D. The 2-pixel error rate is used for Middlebury 2014, and 1-pixel error rate for ETH3D. Bold: Best.
@@ VII More results of UDR
In DR-Stereo, we update the initial disparity map using the UDR module, which efficiently corrects for obvious disparity errors. Table VIII shows the quantitative results.
TABLE VIII: Quantitative results for a single UDR module.We record the results of the UDR updating of the initial disparity on multiple datasets.
@@ IX Other attempts of DR loss
In this section, we explore alternative forms of DR loss. DRloss is a loss function that focuses on small disparity updates. Table IX shows the many forms we have experimented with. We observe that its specific forms can be varied. While the sigmoid form of DR loss achieves more performance gains, we find experimentally that it is difficult to train when the weight bias is small. Therefore, we did not adopt this setting in the end. We will investigate more possibilities of DR loss in the future.
TABLE IX: Other attempts of DR loss. The baseline is IGEV-Stereo.

### v4

@@ Abstract
The updated version will be released soon!!!
@@ V-D Ablation Study
TABLE III: Ablation study and complexity of DR-stereo. The baseline is IGEV-Stereo. The last two columns are the results when the size of input image is $1248\times 384$. Bold: Best.
@@
Combinations of proposed methods: As shown in Table III, we experiment with multiple combinations of the proposed methods. Comparing the results with those in Tables LABEL:tab:UDCUDR and VII, our methods achieve further improvements after combination. In addition, we record the increase in the number of parameters, time consumption after inserting the design modules. It can be seen that inserting the UDR and UDC puts little burden for the model. As we present in Section IV-B, the combination of UDR and UDC does not increase the number of parameters compared to a single method, since they share parameters from the UEC module.
@@ V-F Benchmarks
In this section, we compare DR-Stereo with the state-of-the-art methods published on Scene Flow and KITTI. Tables IV shows the quantitative results. With similar training strategies, DR-Stereo achieves a new SOTA EPE on the Scene Flow test set. Evaluation results on the KITTI benchmark show that DR-Stereo achieves the best performance on the vast majority of metrics. At the time of writing, our method outperforms all published methods on the online KITTI 2015 leaderboard.
TABLE IV: Quantitative evaluation on Scene Flow and KITTI 2015. Bold: Best.
@@ V-G Zero-shot Generalization
We pre-train DR-Stereo on Scene Flow and test it directly on Middlebury 2014 and ETH3D. As shown in Table V, Our DR-Stereo achieves very competitive generalisation performance. Compared with the original optimal method IGEV-Stereo, our method achieves an overall improvement.
TABLE V: Synthetic data generalization experiments. We pre-train our model on Scene Flow and test it directly on Middlebury 2014 and ETH3D. The 2-pixel error rate is used for Middlebury 2014, and 1-pixel error rate for ETH3D. Bold: Best.
@@ VII More results of UDR
In DR-Stereo, we update the initial disparity map using the UDR module, which efficiently corrects for obvious disparity errors. Table VI shows the quantitative results.
TABLE VI: Quantitative results for a single UDR module.We record the results of the UDR updating of the initial disparity on multiple datasets.
@@ IX Other attempts of DR loss
In this section, we explore alternative forms of DR loss. DRloss is a loss function that focuses on small disparity updates. Table VII shows the many forms we have experimented with. We observe that its specific forms can be varied. While the sigmoid form of DR loss achieves more performance gains, we find experimentally that it is difficult to train when the weight bias is small. Therefore, we did not adopt this setting in the end. We will investigate more possibilities of DR loss in the future.
TABLE VII: Other attempts of DR loss. The baseline is IGEV-Stereo.